OpenStudy (hothotleg):

I have 2 questions! I need them answered and explained! I will fan and medal!

OpenStudy (hothotleg):

@DangerousJesse

OpenStudy (anonymous):

What are your questions?

OpenStudy (hothotleg):

Alright, The first one, is this: A new spiral notebook contains 30 more sheets of paper than a new memo book. The total number of sheets of paper in 3 new spiral notebooks and 5 new memo books is 810. Write a system of equations can be used to find s, the number of sheets of paper in one new spiral notebook, and m, the number of sheets of paper in one new memo book.

OpenStudy (hothotleg):

I am completely stumped.

OpenStudy (solomonzelman):

s = m + 30 3s + 5m = 810

OpenStudy (solomonzelman):

(where s - spiral notebook, m - memo book)

OpenStudy (hothotleg):

Oh! Thank you! Is that the explanation? : O

OpenStudy (solomonzelman):

I read the problem and set the equations as I read.

OpenStudy (solomonzelman):

I know that: `spiral notebook contains 30 more sheets of paper than memo book. ` OR, `"s" is 30 more than "m"` And that means that you would have to add 30 (papers) to "m" to make it equivalent to "s". And this can be written as: s=m+30

OpenStudy (solomonzelman):

That is my first equation (s=m+30)

OpenStudy (solomonzelman):

`The number of papers in 3 spiral notebooks and 5 memo books is 810.` The number of papers in one spiral notebook is "s" So, the number of papers in 3 spiral notebooks is "3s" The number of papers in one memo book is "m" So, the number of papers in 5 memo books is "5m" So, you therefore set, `3s + 5m = 810`

OpenStudy (solomonzelman):

and thus, I have obtained my system of equation s=m+30 3s + 5m = 810

OpenStudy (hothotleg):

Oh! That makes sense! Thank you very much! I have one more question, and then I'll be off ^u^

OpenStudy (hothotleg):

At a fundraiser breakfast, the bill for three glasses of orange juice and five pancake specials is $7.60, whereas the bill for one glass of orange juice and two pancake specials is $2.90. The system of equation { 3𝑗 +5𝑝 = 7.60 𝑗 +2𝑝 = 2.90 represents this situation, where j represents the cost of a glass of orange juice and p represents the cost of a pancake special. What would be the bill for one glass of orange juice and one pancake special? Show all work carefully!!

OpenStudy (solomonzelman):

What are the full instructions, to write the system of equation to the problem, or to also solve the system after writing it?

OpenStudy (hothotleg):

Well, I would go with solving it after writing it, to be safe. I want to thank you again for helping me! ^u^

OpenStudy (solomonzelman):

your system for this (last) problem, is correct. Just letting you know and confirming...

OpenStudy (solomonzelman):

One way I would solve the system #1 is...

OpenStudy (solomonzelman):

s = m + 30 3s + 5m = 810 s - m = 30 3s + 5m = 810 5s - 5m = 150 3s + 5m = 810 -------------- 8s + 0 = 960 8s=960 s=120

OpenStudy (solomonzelman):

and then use that to solve for s.

OpenStudy (hothotleg):

AH! Thank you so much!